OpenStudy (df001):

What is the amplitude of y=cos1/2x ?

OpenStudy (anonymous):

y=Acos(Bx-C), A=amplitude, Period=2π/B, Phase shift=C/B

OpenStudy (df001):

But there's no a so is it 1 or 0

OpenStudy (anonymous):

1

OpenStudy (df001):

Ty

OpenStudy (anonymous):

YW